Consider the reaction 3CaCl2(aq) + 2Na3PO4(aq) --> Ca3(PO4)2(s) + 6 NaCl(aq) (molar mass of calcium phosphate is 310.18 g/mol)

What mass of Ca3(PO4)2 is produced from 94.2 mL of 0.250 M Na3PO4 solution and 275 mL of 0.150 M CaCl2 solution? _____ g (Give answer to correct sig figs in units of g)
1
Expert's answer
2020-04-03T11:48:40-0400

Here the given reaction is ,


"3CaCl_2(aq) +2Na_3PO_4(aq)\longrightarrow Ca_3(PO_4)_2(s)+6 NaCl"


Here amount of "Na_3PO_4" is 94.2 mL of 0.25M solution


moles of Na3PO4= "Molarity\times volume (in L)=0.0942\times0.25=0.02355" mole


similarly , moles of "CaCl_2= molarity\times volume= 0.15\times 0.275= 0.04125" mole


from above reaction we can say that, 3 mole of CaCl2 react with 2 mole of Na3PO4


0.04125 mole of CaCl2 react with = "\frac{2}{3} \times 0.04125=0.0275" mole so from here we are


getting the limiting reagent as Na3PO4


so from the above reaction we can say that ,

2 mole of "Na_3PO_4" gives 1 mole of "Ca_3(PO_4)_2"


so, 0.02355 mole of "Na_3PO_4" gives = "\frac{0.02355}{2}= 0.011775" mole of "Ca_3(PO_4)_2"


so mass of "Ca_3(PO_4)_2= mole\times molar mass= 0.011775\times 310= 3.6425" g